# Asymptotics of spherical dynamos exhibiting a small-scale MAC balance

### v1

@@
Asymptotics of spherical dynamos exhibiting a small-scale MAC balance
Justin A. Nicoski1    Andy Esseln1    Chris Davies2 and Michael A. Calkins1
1Department of Physics
   University of Colorado    Boulder    Colorado 80309    USA
2School of Earth and Environment
   University of Leeds    Leeds LS2 9JT    United Kingdom
keywords:
Dynamo: theories and simulations; Core; Numerical modelling
{summary}
Understanding the asymptotic behaviour of numerical dynamo models is critical for extrapolating results to the physical conditions that characterise terrestrial planetary cores. Here we investigate the behaviour of convection-driven dynamos reaching a MAC (magnetic-Archimedes-Coriolis) balance on the convective length scale and compare the results with non-magnetic convection cases. In particular, the dependence of physical quantities on the Ekman number, $Ek$, is studied in detail. The scaling of velocity dependent quantities is observed to be independent of the force balance and in agreement with quasi-geostrophic theory. The primary difference between dynamo and non-magnetic cases is that the fluctuating temperature is order unity in the former such that the buoyancy force scales with the Coriolis force. The MAC state yields a scaling for the flow speeds that is identical to the so-called CIA (Coriolis-inertia-Archimedes) scaling. There is an $O(Ek^{1/3})$ length scale present within the velocity field irrespective of the leading order force balance. This length scale is consistent with the asymptotic scaling of the terms of the governing equations and is not an indication that viscosity plays a dominant role. The peak of the kinetic energy spectrum and the ohmic dissipation length scale both exhibit an Ekman number dependence of approximately $Ek^{1/6}$, which is consistent with a scaling of $Rm^{-1/2}$, where $Rm$ is the magnetic Reynolds number. For the dynamos, advection remains comparable to, and scales similarly with, both inertia and viscosity, implying that nonlinear convective Rossby waves play an important role in the dynamics even in a MAC regime.
@@ 1 Introduction
The magnetic fields of the Earth, planets and stars are all thought to be generated by the motion of electrically conducting fluids (e.g. Jones, 2011). In the case of the Earth, this dynamo process occurs in the liquid iron outer core (Roberts & King, 2013). The core is likely a turbulent environment and therefore characterised by a broad range of spatial and temporal scales (Aurnou et al., 2015), which has necessitated the use of numerical models for studying its dynamics (e.g. Zhang & Busse, 1989; Kageyama et al., 1995; Glatzmaier & Roberts, 1995). These early simulations exhibited magnetic fields that show many similarities with the geomagnetic field, including a predominantly dipolar structure and the presence of polarity reversals (Glatzmaier & Roberts, 1995). Since Glatzmaier & Roberts (1995), many dynamo simulations have been performed which have qualitative similarities with the Earth’s magnetic field (e.g. Christensen et al., 2010; Meduri et al., 2021). As such, dynamo simulations are believed to capture certain aspects of the dynamics of the geodynamo.
@@
Due to their high computational cost, direct numerical simulations (DNS) cannot yet reach the parameter space of the geodynamo. In particular, the Ekman number, $Ek=\nu/(\Omega H^{2})$, characterising the ratio of the viscous force to the Coriolis force at large length scales, is approximately $Ek\sim 10^{-15}$ in the outer core (here $\nu$ is the kinematic viscosity, $\Omega$ is the angular frequency or rotation rate and $H$ is the depth of the fluid layer). In contrast, DNS of spherical dynamos rarely reach Ekman numbers smaller than $Ek\sim 10^{-7}$ (Schaeffer et al., 2017). Another important nondimensional parameter is the magnetic Prandtl number, which is defined as $Pm=\nu/\eta$, where $\eta$ is the magnetic diffusivity. Studies indicate that $Pm=O(10^{-6})$ in the Earth (e.g. Pozzo et al., 2013), which is also far beyond what can currently be investigated in DNS due to the large flow speeds necessary to sustain dynamo action in such a regime.
@@
There are three modeling strategies that can be used to help bridge this gap. Historically, the primary approach consists of carrying out DNS at ever more extreme parameters and developing scaling laws from the output such that predictions can be made in an Earth-like regime (e.g. Christensen & Aubert, 2006; Davidson, 2013). Scaling laws exploit dominant balances in the governing equations and must make assumptions with regard to how the spatiotemporal dynamics depend on system parameters. Another approach involves assuming that the small-scales are of negligible dynamical importance such that only the large scales in the system are directly modeled. Hyperviscous damping on the smallest simulated scales can be used for numerical stabilisation, allowing for considerable extension of the parameter space (Aubert et al., 2017; Guervilly et al., 2019; Aubert, 2023). A third approach is a model reduction strategy that involves identifying the small parameters of the system and using perturbation methods to formally derive a new, simpler, equation set that is significantly cheaper to solve, thereby allowing access to more extreme parameter space. This strategy has been successful in the planar (Julien et al., 1998; Sprague et al., 2006; Julien et al., 2012; Calkins et al., 2015) and ‘annulus’ models (Busse, 1970; Calkins et al., 2013) but has not been applied to the nonlinear spherical problem due in part to the variety of length scales present. Towards this end, one of the main goals of the present work is to quantify the Ekman number dependence of the small-scale (i.e. convection-scale) balances in the spherical geometry when strong magnetic field is present, which is a crucial aspect for understanding the asymptotic scaling of the system.
@@
The derivation of scaling laws and asymptotic expansions relies on identification of the dominant force balance in the dynamics. Numerical simulations of rapidly rotating dynamos have revealed two main force balances, depending on the relative size of the Lorentz force, which is controlled primarily by the value of $Pm$. When the Lorentz force is sufficiently small, the leading order force balance is between the Coriolis force and the pressure gradient force and the resulting dynamics are quasi-geostrophic (QG) (Soderlund et al., 2012; Yadav et al., 2016; Calkins, 2018; Yan & Calkins, 2022a, b). Some studies find that QG dynamos exhibit many quantitative similarities with non-magnetic QG convection (Soderlund et al., 2012; Yan & Calkins, 2022b). In this sense, these simulations are similar to rapidly rotating convective cases without a magnetic field that develop a balance between the Coriolis and pressure forces at zeroth order, with the rest of the terms entering at first order, as described in asymptotic models (e.g. Sprague et al., 2006). Based on this line of reasoning, asymptotic dynamo models in planar geometries can be derived assuming that the Lorentz force enters at the same order as the first-order terms that appear in asymptotic quasi-geostrophic convection models (such as the viscous force, the advective term, and the buoyancy force) (Calkins et al., 2015). Simulations of rotating plane layer dynamos at small magnetic Prandtl numbers find the predicted asymptotic scalings from the quasi-geostrophic dynamo model, with the Lorentz force entering at the same order as the viscous force (Yan & Calkins, 2022a). Even within the QG limit, however, it is possible to generate magnetic field strengths such that the dynamics are vastly different relative to non-magnetic rotating convection (e.g. Plumley et al., 2018; Maffei et al., 2019).
@@
In contrast to the geostrophic balance, a MAC (magnetic-Archimedes-Coriolis) balance occurs when the Lorentz force is comparable in magnitude to the Coriolis and buoyancy forces (Roberts & King, 2013). It has long been suggested that the Earth might be in a such a regime (Taylor, 1963), and numerical simulations also suggest this possibility can arise on certain length scales (e.g. Soderlund et al., 2015; Dormy, 2016; Aurnou & King, 2017). Numerical simulations can only achieve this regime with unrealistically large values of $Pm$ (e.g. Dormy, 2016; Petitdemange, 2018; Menu et al., 2020). However, such studies are nevertheless important for understanding the differences between QG and MAC dynamos. On the other hand, scale-dependent force balances often suggest that the primary force balance is QG at large length scales, so it is sometimes argued that dynamos are QG-MAC rather than simply MAC, at least at large length scales (Aubert et al., 2017; Schwaiger et al., 2019, 2021).
@@
Given that the force balance can be very different for dynamo simulations with a strong magnetic field as compared to convective cases without a magnetic field, the question arises as to how the magnetic field might alter the asymptotic relations found in rotating convection. We study this question by comparing the dynamo cases from Calkins et al. (2021) with non-magnetic rotating convection cases and comparing the asymptotic behaviour of the two sets of data. The dynamo cases from Calkins et al. (2021) are run at a fixed $Pm=2$, and develop a strong magnetic field, especially as the Ekman number is reduced. For many of our cases, the fluctuating Lorentz force is over half as large as the fluctuating Coriolis force, which ensures that our dynamo simulations are strongly influenced by the magnetic field. The paper is structured as follows. In section 2, the model equations, numerical method, and outputs are described. Section 3 briefly summarizes the expected asymptotic scalings of flow speeds, length scales, and forces in non-magnetic rapidly rotating convection. Section 4 compares the asymptotic scalings between the non-magnetic and dynamo cases. Finally, a discussion of the results is provided in section 5.
@@ 2 Model
Here, $Ra$ is the Rayleigh number, $Pr$ is the Prandtl number, $g_{o}$ is the acceleration due to gravity at the outer boundary, $\alpha$ is the thermal expansion coefficient, $\kappa$ is the thermal diffusivity, and $\eta$ is the magnetic diffusivity. For our simulations, we fix $Pr=1$ and most of the dynamo simulations use $Pm=2$. A subset of simulations are carried out for variable magnetic Prandtl number, ranging from $Pm=0.2$ up to $Pm=5$. We employ no-slip boundary conditions for the velocity field and electrically insulating boundary conditions for the magnetic field. The non-magnetic cases are identical to the dynamo cases except that $\mathbf{B}=0$.
We use the pseudo-spectral code Rayleigh to numerically solve the governing equations (Featherstone et al., 2022). Rayleigh uses spherical harmonics to represent functions over shells, and Chebyshev polynomials to represent functions in radius. A full $3/2$ de-aliasing is used for both the spherical harmonics and the Chebyshev polynomials. Time stepping is performed with a 2nd-order Crank-Nicolson/Adams-Bashforth scheme. Rayleigh has been successfully benchmarked against the cases in Christensen et al. (2001), and has been used in many previous studies (e.g. Featherstone & Hindman, 2016; Calkins et al., 2021; Heimpel et al., 2022; Nicoski et al., 2024).
@@ 2.1 Notation and outputs
where the prime denotes that the $m=0$ mode has been excluded in order to remove the influence of large-scale flows and fields. $l_{peak}$ is defined as the degree at which $E^{\prime}_{l}$ achieves a maximum. To calculate $l_{peak}$, we use a simple polynomial interpolation of $E^{\prime}_{l}$ before finding the peak, as other studies have done (Guervilly et al., 2019). A peak length scale can then be defined as
@@
respectively, where $\boldsymbol{\omega}=\nabla\times\mathbf{u}$ is the vorticity. The Taylor microscales are generally considered to be representative of the length scales at which diffusion (momentum/magnetic field) becomes important. For this reason, they are often referred to as dissipation length scales. Length scales are also computed directly from various terms in the momentum equation. We define the length scale computed from the advection term in the momentum equation as
@@
We remove the $\ell=0$ part of the nonlinear terms since the $\ell=0$ component is not dynamically relevant in the radial component of the momentum equation for an incompressible fluid. Except for the buoyancy force which only appears in the radial direction and the viscous force which has significant boundary layers in the $\theta$ and $\phi$ directions, the different components of the forces exhibit identical asymptotic scalings, as can be deduced from the study of Naskar et al. (2025). As also shown in Naskar et al. (2025), the inclusion of Ekman boundary layers in the volume integration can have a significant influence on the observed balances. However, we find that the balances in the radial component of the momentum equation are not sensitive to whether the Ekman boundary layers are included, likely because the velocity normal to the boundary is asymptotically smaller than the horizontal velocity within the boundary layer (e.g. Greenspan, 1968).
In the limit $Ek\rightarrow 0$, it is well known that the critical Rayleigh number required to drive convection scales as $Ra_{c}=O(Ek^{-4/3})$ (Roberts, 1968). Thus, the key control parameter in rapidly rotating convection-driven dynamos is the asymptotically reduced Rayleigh number defined as (Julien et al., 1998)
@@ 3 Quasi-geostrophic scalings
Here we briefly outline the known scalings from QG theory, which applies when the Coriolis force and the pressure gradient force enter at leading order, whereas all remaining terms in the momentum equation, including the Lorentz force, enter the dynamics at the next order (e.g. Calkins, 2018). A simple derivation of the QG scalings in a plane layer can be derived by assuming that all of the terms in the vorticity equation enter at the same asymptotic order in Ekman number. It is also assumed that length scales along the direction of the rotation axis are order one, and the length scales perpendicular to the rotation axis are all of the same asymptotic order. This implies
$O\left(Ek^{-1}Re\right)=O\left(Re^{2}\ell^{-2}\right)=O\left(Re\ell^{-3}\right)=O\left(RaT^{\prime}\ell^{-1}\right),$ (18)
where order one terms have been dropped and $\ell$ is the length scale. This system of equations implies implies that the Reynolds number ($Re=uH/\nu$) scales as $O(Ek^{-1/3})$, the fluctuating temperature scales as $O(Ek^{1/3})$, and the horizontal length scale scales as $O(Ek^{1/3})$. These are the same asymptotic scalings that are used in reduced models (Sprague et al., 2006). Note that these scalings only give the dependence on $Ek$, but not on $\widetilde{Ra}$. In our non-dimensionalisation, the Coriolis force and pressure force are then expected to scale as $O(Ek^{-4/3})$, while the viscous force, buoyancy force, and advection term are expected to scale as $O(Ek^{-1})$. We will test our data against these predicted QG scalings, though it is not obvious a priori that the dynamo cases should follow these non-magnetic scalings. Further discussion of QG scalings in spherical convection can be found in Nicoski et al. (2024).

### v3

@@
Abstract Understanding the asymptotic behaviour of numerical dynamo models is critical for extrapolating results to the physical conditions that characterise terrestrial planetary cores. Here we investigate the behaviour of convection-driven dynamos reaching a MAC (magnetic-Archimedes-Coriolis) balance on the convective length scale and compare the results with non-magnetic convection cases. In particular, the dependence of physical quantities on the Ekman number, $Ek$, is studied in detail. The scaling of velocity dependent quantities is observed to be independent of the force balance and in agreement with quasi-geostrophic theory. The primary difference between dynamo and non-magnetic cases is that the fluctuating temperature is order unity in the former such that the buoyancy force scales with the Coriolis force. The MAC state yields a scaling for the flow speeds that is identical to the so-called CIA (Coriolis-inertia-Archimedes) scaling. There is an $O(Ek^{1/3})$ length scale present within the velocity field irrespective of the leading order force balance. This length scale is consistent with the asymptotic scaling of the terms of the governing equations and is not an indication that viscosity plays a dominant role. The peak of the kinetic energy spectrum and the ohmic dissipation length scale both exhibit an Ekman number dependence of approximately $Ek^{1/6}$, which is consistent with a scaling of $Rm^{-1/2}$, where $Rm$ is the magnetic Reynolds number. For the dynamos, advection remains comparable to, and scales similarly with, both inertia and viscosity, implying that nonlinear convective Rossby waves play an important role in the dynamics even in a MAC regime.
@@ 1 Introduction
The magnetic fields of the Earth, planets and stars are all thought to be generated by the motion of electrically conducting fluids [cJ11b, e.g.]. In the case of the Earth, this dynamo process occurs in the liquid iron outer core [pR13]. The core is likely a turbulent environment and therefore characterised by a broad range of spatial and temporal scales [jmA15], which has necessitated the use of numerical models for studying its dynamics [kZ89, aK95, gG95, e.g.]. These early simulations exhibited magnetic fields that show many similarities with the geomagnetic field, including a predominantly dipolar structure and the presence of polarity reversals [gG95]. Since [gG95], many dynamo simulations have been performed which have qualitative similarities with the Earth’s magnetic field [uC10b, dM21, e.g.]. As such, dynamo simulations are believed to capture certain aspects of the dynamics of the geodynamo.
@@
Due to their high computational cost, direct numerical simulations (DNS) cannot yet reach the parameter space of the geodynamo. In particular, the Ekman number, $Ek=\nu/(\Omega H^{2})$, characterising the ratio of the viscous force to the Coriolis force at large length scales, is approximately $Ek\sim 10^{-15}$ in the outer core (here $\nu$ is the kinematic viscosity, $\Omega$ is the angular frequency or rotation rate and $H$ is the depth of the fluid layer). In contrast, DNS of spherical dynamos rarely reach Ekman numbers smaller than $Ek\sim 10^{-7}$ [nS17, yL25]. Another important nondimensional parameter is the magnetic Prandtl number, which is defined as $Pm=\nu/\eta$, where $\eta$ is the magnetic diffusivity. Studies indicate that $Pm=O(10^{-6})$ in the Earth [mP13, e.g.], which is also far beyond what can currently be investigated in DNS due to the large flow speeds necessary to sustain dynamo action in such a regime.
@@
There are three modeling strategies that can be used to help bridge this gap. Historically, the primary approach consists of carrying out DNS at ever more extreme parameters and developing scaling laws from the output such that predictions can be made in an Earth-like regime [uC06, pD13b, e.g.]. Scaling laws exploit dominant balances in the governing equations and must make assumptions with regard to how the spatiotemporal dynamics depend on system parameters. Another approach involves assuming that the small-scales are of negligible dynamical importance such that only the large scales in the system are directly modeled. Hyperviscous damping on the smallest simulated scales can be used for numerical stabilisation, allowing for considerable extension of the parameter space [jA17, cG19, jA23]. A third approach is a model reduction strategy that involves identifying the small parameters of the system and using perturbation methods to formally derive a new, simpler, equation set that is significantly cheaper to solve, thereby allowing access to more extreme parameter space. This strategy has been successful in the planar [kJ98a, mS06, kJ12, mC15b] and ‘annulus’ models [fB70, mC13] but has not been applied to the nonlinear spherical problem due in part to the variety of length scales present. Towards this end, one of the main goals of the present work is to quantify the Ekman number dependence of the small-scale (i.e. convection-scale) balances in the spherical geometry when strong magnetic field is present, which is a crucial aspect for understanding the asymptotic scaling of the system.
@@
The derivation of scaling laws and asymptotic expansions relies on identification of the dominant force balance in the dynamics. Numerical simulations of rapidly rotating dynamos have revealed two main force balances, depending on the relative size of the Lorentz force, which is controlled primarily by the value of $Pm$. When the Lorentz force is sufficiently small, the leading order force balance is between the Coriolis force and the pressure gradient force and the resulting dynamics, which arise from perturbations to this balance, are termed quasi-geostrophic (QG) [kS12, rY16, mC18, mY22, mY22b]. Some studies find that QG dynamos exhibit many quantitative similarities with non-magnetic QG convection [kS12, mY22b]. In this sense, these simulations are similar to rapidly rotating convective cases without a magnetic field that develop a balance between the Coriolis and pressure forces at zeroth order, with the rest of the terms entering at first order, as described in asymptotic models [mS06, e.g.]. Based on this line of reasoning, asymptotic dynamo models in planar geometries can be derived assuming that the Lorentz force enters at the same order as the first-order terms that appear in asymptotic quasi-geostrophic convection models (such as the viscous force, the advective term, and the buoyancy force) [mC15b]. Simulations of rotating plane layer dynamos at small magnetic Prandtl numbers find the predicted asymptotic scalings from the quasi-geostrophic dynamo model, with the Lorentz force entering at the same order as the viscous force [mY22]. Even within the QG limit, however, it is possible to generate magnetic field strengths such that the dynamics are vastly different relative to non-magnetic rotating convection [mP18, sM19, e.g.].
@@
In contrast to the geostrophic balance, a MAC (magnetic-Archimedes-Coriolis) balance occurs when the Lorentz force is comparable in magnitude to the Coriolis and buoyancy forces [pR13]. It has long been suggested that the Earth might be in a such a regime [jT63], and numerical simulations also suggest this possibility can arise on certain length scales [kS15, eD16, jmA17, e.g.]. Numerical simulations can only achieve this regime with unrealistically large values of $Pm$ [eD16, lP18, mM20, e.g.]. However, such studies are nevertheless important for understanding the differences between QG and MAC dynamos. On the other hand, scale-dependent force balances often suggest that the primary force balance is QG at large length scales, so it is sometimes argued that dynamos are QG-MAC rather than simply MAC, at least at large length scales [jA17, tS19, tS21].
@@
Given that the force balance can be very different for dynamo simulations with a strong magnetic field as compared to convective cases without a magnetic field, the question arises as to how the magnetic field might alter the asymptotic relations found in rotating convection. We study this question by comparing the dynamo cases from [mC21] with non-magnetic rotating convection cases and comparing the asymptotic behaviour of the two sets of data. The dynamo cases from [mC21] are run at a fixed $Pm=2$, and develop a strong magnetic field, especially as the Ekman number is reduced. For many of our cases, the fluctuating Lorentz force is over half as large as the fluctuating Coriolis force, which ensures that our dynamo simulations are strongly influenced by the magnetic field. The paper is structured as follows. In section 2, the model equations, numerical method, and outputs are described. Section 3 briefly summarizes the expected asymptotic scalings of flow speeds, length scales, and forces in non-magnetic rapidly rotating convection. Section 4 compares the asymptotic scalings between the non-magnetic and dynamo cases. Finally, a discussion of the results is provided in section 5.
@@ 2 Model
Here, $Ra$ is the Rayleigh number, $Pr$ is the Prandtl number, $g_{o}$ is the acceleration due to gravity at the outer boundary, $\alpha$ is the thermal expansion coefficient, $\kappa$ is the thermal diffusivity, and $\eta$ is the magnetic diffusivity. For our simulations, we fix $Pr=1$ and most of the dynamo simulations use $Pm=2$. A subset of simulations are carried out for variable magnetic Prandtl number, ranging from $Pm=0.2$ up to $Pm=5$. Four different values of the Ekman number are used in the present study: $Ek=(3\times 10^{-6},10^{-5},3\times 10^{-5},10^{-4})$. The Rayleigh number is varied up to approximately 50 times the critical Rayleigh number for the onset of non-magnetic convection. We employ no-slip boundary conditions for the velocity field and electrically insulating boundary conditions for the magnetic field. The non-magnetic cases are identical to the dynamo cases except that $\mathbf{B}=0$.
We use the pseudo-spectral code Rayleigh to numerically solve the governing equations [featherstone2022]. Rayleigh uses spherical harmonics to represent functions over shells, and Chebyshev polynomials to represent functions in radius. A full $3/2$ de-aliasing is used for both the spherical harmonics and the Chebyshev polynomials. Time stepping is performed with a 2nd-order Crank-Nicolson/Adams-Bashforth scheme. Rayleigh has been successfully benchmarked against the cases in [uC01b], and has been used in many previous studies [nF16, mC21, mH22, jN24, e.g.].
@@ 2.1 Notation and outputs
where the prime denotes that the $m=0$ mode has been excluded in order to remove the influence of large-scale flows and fields. $l_{peak}$ is defined as the degree at which $E^{\prime}_{l}$ achieves a maximum. To calculate $l_{peak}$, we use a simple polynomial interpolation of $E^{\prime}_{l}$ before finding the peak, as other studies have done [cG19]. A peak length scale can then be defined as
@@
respectively, where $\bm{\omega}=\nabla\times\mathbf{u}$ is the vorticity. The Taylor microscales are generally considered to be representative of the length scales at which diffusion (momentum/magnetic field) becomes important. For this reason, they are often referred to as dissipation length scales. Length scales are also computed directly from various terms in the momentum equation. We define the length scale computed from the advection term in the momentum equation as
@@
We remove the $\ell=0$ part of the nonlinear terms since the $\ell=0$ component is not dynamically relevant in the radial component of the momentum equation for an incompressible fluid. Except for the buoyancy force which only appears in the radial direction and the viscous force which has significant boundary layers in the $\theta$ and $\phi$ directions, the different components of the forces exhibit identical asymptotic scalings, as can be deduced from the study of [sN25]. As also shown in [sN25], the inclusion of Ekman boundary layers in the volume integration can have a significant influence on the observed balances. However, we find that the balances in the radial component of the momentum equation are not sensitive to whether the Ekman boundary layers are included, likely because the velocity normal to the boundary is asymptotically smaller than the horizontal velocity within the boundary layer [hG68, e.g.].
In the limit $Ek\rightarrow 0$, it is well known that the critical Rayleigh number required to drive convection scales as $Ra_{c}=O(Ek^{-4/3})$ [pR68]. Thus, the key control parameter in rapidly rotating convection-driven dynamos is the asymptotically reduced Rayleigh number defined as [kJ98a]
@@ 3 Quasi-geostrophic scalings
Here we briefly outline the known scalings from QG theory, which applies when the Coriolis force and the pressure gradient force enter at leading order, whereas all remaining terms in the momentum equation, including the Lorentz force, enter the dynamics at the next order [mC18, e.g.]. A simple derivation of the QG scalings in a plane layer can be derived by assuming that all of the terms in the fluctuating vorticity equation enter at the same asymptotic order in Ekman number. It is also assumed that length scales along the direction of the rotation axis are order one, and the length scales perpendicular to the rotation axis are all of the same asymptotic order. We note that multiple length scales are present in the radial direction in spherical geometries, though the larger length scale exhibits a weaker asymptotic dependence on the Ekman number [cJ00, eD04, e.g.]; we discuss this point further in the results. Curling inertia and advection give respectively,
$\nabla\times\left({\partial_{t}}\mathbf{u}^{\prime}\right)=O\left(\tau^{-1}\ell^{-1}Re_{c}\right),\quad\nabla\times\left(\mathbf{u}^{\prime}\cdot\nabla\mathbf{u}^{\prime}\right)=O\left(Re_{c}^{2}\ell^{-2}\right),$ (18)
where $\tau$ is a characteristic timescale, $\ell$ is the lengthscale and $\mathbf{u}^{\prime}=O(Re_{c})$. We have ignored the $\phi$-averaged component of the advection term since it is smaller than the term included above. For inertia and advection to be asymptotically comparable, we require the timescale to be
$\tau=O\left(\ell Re_{c}^{-1}\right).$ (19)
Curling the viscous and buoyancy forces gives, respectively,
$\nabla\times\left(\nabla^{2}\mathbf{u}\right)=O\left(\ell^{-3}Re_{c}\right),$ (20)
$\nabla\times\left(\frac{Ra}{Pr}\frac{r}{r_{o}}T^{\prime}\,{\bf\widehat{r}}\right)=O\left(\ell^{-1}RaT^{\prime}\right),$ (21)
where order one terms have been dropped. The above system of equations implies the following scalings:
$Re=O\left(E^{-1/3}\right),\quad T^{\prime}=O\left(E^{1/3}\right),$ (22)
$\ell=O\left(E^{1/3}\right),\quad\tau=O\left(E^{-2/3}\right).$ (23)
The above scalings are identical to those used in reduced models of rapidly rotating convection and rotating convection-driven dynamos [mS06, mC15b]. Importantly, these scalings arise from the dominant (leading order) geostrophic force balance on the small $O(Ek^{1/3})$ horizontal length scale, and the theory is not limited to weakly nonlinear, or weakly forced, systems. The presence of $O(Ek^{1/3})$ in the above scalings does not arise from the assumption that the viscous force is dominant in the vorticity equation since the same dependence can arise without reference to viscosity. Indeed, $\widetilde{Ra}$ now controls the relative sizes of these terms; this point is addressed in the results.
In our non-dimensionalisation, the Coriolis force and pressure force are then expected to scale as $O(Ek^{-4/3})$, while the viscous force, buoyancy force, and advection term are expected to scale as $O(Ek^{-1})$. These QG scalings are tested with the dynamo simulation data, though it is not obvious a priori that the dynamo cases should follow these non-magnetic scalings when the Lorentz force becomes comparable to the Coriolis force. Further discussion of QG scalings in spherical convection can be found in [jN24].